Given that n is a positive integer and in the expansion of (1+ax)n, the coefficient of x is 200. Find the coefficient of x in the expansion of (1+ax)2n.


1
Expert's answer
2021-10-27T15:25:04-0400

"(1+ax)^n:"


"\dbinom{n}{1}(ax)(1)^{n-1}=nax=200x"

"na=200"

"(1+ax)^{2n}:"


"\dbinom{2n}{1}(ax)(1)^{2n-1}=2nax=2(200)x=400x"

The coefficient of "x" in the expansion of "(1+ax)^{2n}" is "400."